OpenStudy (anonymous):

.What is the volume of a cone with a radius of 6 cm and a height of 12 cm? Use 3.14 for pi. Round your answer to the nearest hundredth. 452.16 cm3 465.24 cm3 472.12 cm3 480.50 cm3

OpenStudy (anonymous):

\[Volume ~of~the~cone=\frac{ 1 }{ 3 }\pi r^2 h\] r=radius of the cone h=height of the cone